OpenStudy (anonymous):

Can someone check my answer? Medal! 3. divide x^2+2x+1/x^2-1 divided by (2x^2+2x) a. 1/2x(x-1) - my answer b. x+1/2x c. 1/2 d. 2x(x+1)^2/x-1

OpenStudy (anonymous):

\[x^2+2x+1/x^2-1 \div (2x^2+2)\]

OpenStudy (tkhunny):

You ALMOST made it better. Not quite. Do you mean \([(x^{2} +2x+1)/(x^{2} −1)]/(2x^{2} +2) \)

OpenStudy (anonymous):

this is it

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

Why don't you write that? Parentheses are not necessary in the typeset version. The parentheses are ABSOLUTELY necessary for th ein-line version. Make SURE it says what you want. Okay, how did you proceed? Factor everything, i presume?

OpenStudy (anonymous):

yes thats how I started it, and i'm trying to learn how to do equations like these so I am still trying to learn the steps

OpenStudy (tkhunny):

The first step is to learn how to write it. You MUST see why the parentheses are not optional. Second, you need to see that this is NOT an "equation". It is an algebraic expression. Equations have one of these "=" in them. \(x^{2}+2x+1 = (x+1)^{2}\) \(x^{2}-1 = (x+1)(x-1)\) \(2x^{2} + 2x = 2x(x+1)\) Thus: \(\dfrac{\dfrac{x^{2}+2x+1}{x^{2}-1}}{2x^{2} + 2x} = \dfrac{(x+1)^{2}}{(x+1)(x-1)\cdot 2x(x+1)} = \dfrac{1}{(x-1)\cdot 2x}\) It looks like you have it. Why do you doubt?

OpenStudy (anonymous):

Im just not very good at math and it takes me awhile to learn these things lol

OpenStudy (tkhunny):

Hogwash. If you sifted through all that, you are now, officially, no longer "not very good at math". It is time for you to stop saying that and stop thinking that.

OpenStudy (anonymous):

Thank you!

OpenStudy (tkhunny):

Please note that you said what your answer was but you provided no evidence that it was YOUR answer. As far as this thread knows, you could have copied the answer out of a book. You MUST show your work.